OpenStudy (anonymous):

implicit derivative of y^2-xy-3x=1

OpenStudy (tkhunny):

Well, do it.

OpenStudy (anonymous):

i think that it's 2yy'-x'y* -xy' -3x=0

OpenStudy (anonymous):

is that right?

OpenStudy (tkhunny):

Not bad. One piece at a time. \(\dfrac{d}{dx}y^{2} = 2yy'\) \(\dfrac{d}{dx}-xy = -(xy' + y)\) ==> x' = 1 :-) \(\dfrac{d}{dx}-3x = -3\) ==> You didn't do anything to this piece. \(\dfrac{d}{dx}1 = 0\)

OpenStudy (anonymous):

i forgot product rule is f'g+fg'

OpenStudy (anonymous):

i thought it was f'g * fg'

OpenStudy (anonymous):

why do we take derivative of 3x though? I thought we were just taking the derivative of y

OpenStudy (anonymous):

so then y'=-1

OpenStudy (anonymous):

thanks a ton

OpenStudy (tkhunny):

Have to get them all. Good work. 2yy' - xy' - y - 3 = 0 y'(2y - x) = y+3 \(y' = \dfrac{y+3}{2y-x}\) How did you get -1 out of that?